OpenStudy (anonymous):

A solution of salt and water contains 70 grams of water per 140 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 35 milliliters of the solution. [Round off your answer to two decimal places] @TheRealMeeeee

OpenStudy (anonymous):

@adrynicoleb

OpenStudy (anonymous):

@tester97

OpenStudy (therealmeeeee):

We have 70 grams of water/140 mL of solution. 1 mole of water weighs 16 grams. x moles of water would be in 35 mL of solution. Start off by figuring out how grams of water you need for 16 mL of solution. 70 g/140 mL=x g/35 mL x=17.5 g I believe Now that we know we need 8 grams for 16 mL of water, we simply must convert it to moles with some stoichiometry. 17.5 g*(1 mol/16 g)=1.09 moles Notice how I multiplied by what is called the molar mass ratio. The grams canceled out and I am left with moles.

OpenStudy (tester97):

yea wat he said ^^^